# Assurance of AI Systems From a Dependability Perspective

### v2

@@ 1.2 From Assurance to Dependability
In commercial airplanes, “catastrophic failure conditions” (those “which would prevent continued safe flight and landing”) must be “so unlikely that they are not anticipated to occur during the entire operational life of all airplanes of one type” [56]. The “entire operational life of all airplanes of one type” is about $10^{8}$ to $10^{9}$ flights for modern airplanes. With an average flight duration of about 90 minutes, this requires a critical failure rate no worse than about $10^{-9}$ per hour. [^5] There are many measures for reliability, such as failure rate, probability of failure on demand, probability of fatality per mile, etc. And the underlying system models may use discrete or continuous time, with Bernoulli or Poisson failure processes, etc. The general conclusions drawn here are robust to all these choices and so we do not describe them in detail.
@@
Cars are among the most dangerous consumer goods with about 40,000 deaths per year in the United States and a fatal accident rate of a little over 10 per billion miles. It is intended that self-driving cars should be safer than human drivers, so it might seem reasonable (even though driving error is not the only cause of accidents) to set the target at no more than 1 death per billion miles, which is around $10^{-7}$ per hour, given an average speed of 30 mph. However, that is a rather technical assessment that pays no attention to likely public reaction. Herbert Diess, the former CEO of Volkswagen is quoted on their website with a more realistic assessment: “A ratio of ten-to-one is nowhere near good enough. We have approximately 3,200 traffic fatalities in Germany each year. It would be a disaster if we had even 320 deaths due to driverless cars.” Thus, it is widely accepted that the safety target for self-driving cars should be 100 or even 1,000 times better than human drivers, [^6] This sensitivity is illustrated by a recent accident in San Francisco: a car with a human driver hit a pedestrian and threw them under the wheels of a Cruise self-driving taxi, which then dragged the victim for several yards. Subsequently, Cruise lost their license to operate in San Francisco (although their poor initial response to the incident probably contributed to this harsh reaction) [97]. which brings us into, or even beyond, the requirements for commercial aircraft. But commercial aircraft do get assured and certified and their safety record justifies this, so why is $10^{-9}$ seen as such a challenge?
@@
The answer is that we do not assure aircraft solely by testing. If we test a system for $n$ hours and observe no failures, then in the absence of other information, the best prediction we can make is that the likelihood of no failures in another $n$ hours is about 50–50 [113, p. 73, and sidebar on p. 74]. Hence, to secure assurance for failure rates of $10^{-9}$ per hour we would need to test the system for about $10^{9}$ hours, or around 115,000 years. Even with 1,000 copies of the system on test, this is still well over 100 years of continuous operation and is completely infeasible [37, 113, 90].
@@
The purpose of the classical methods of assurance described earlier is to give justified confidence that the system contains no (dependability) faults and hence will have no dependability failures. Confidence can be expressed as a subjective probability, so if we are $95\%$ confident that traditional assurance works, that means we estimate there is only 5% chance that the assured system contains faults. We can now use testing to explore the existence of those potential faults but, unlike the previous case, we know something about the system so when we see $n$ hours with no failures, we can conclude (by what is called Conservative Bayesian Inference, CBI) that we are likely to see another $10n$ with no failures [163, 182]. This reduces the amount of testing required and another idea reduces it still further. This is Bootstrapping [22]: we need assurance for $10^{9}$ hours, but this is over the lifetime of the system. When the system is first deployed, we might be satisfied to know there will be no dependability failures in the first year, and we will have only a few instances of the system operational in that time. So we might need confidence for only, say, $10^{5}$ hours, and testing for this, given prior assurance and CBI, requires only $10^{4}$ hours, which is perfectly feasible. After the first year, we might seek confidence for the next year and for the larger number of systems now installed, but we will have the operational experience of the first year and that should be sufficient (given the tenfold multiplier of CBI) to deliver the required confidence in safe operation, and so on for subsequent years.
@@
But suppose experience in operation does reveal a failure, hopefully not catastrophic (in commercial aircraft, for example, it is required that no catastrophic failure may be caused by a single fault). Any such failure reveals an unanticipated fault, arrival rate, circumstance, or hazard, and these could be precursors to a catastrophic failure. Consequently, commercial airplanes operate in a legal and ethical framework where all incidents and accidents are promptly reported and dispassionately investigated. The FAA issues Airworthiness Directives mandating workarounds or corrections to detected faults; in extreme cases it may temporarily ground the fleet (as it did for the 737 MAX in 2019–2020). Bishop [21] constructs a statistical model for this scenario and shows that, under plausible assumptions, detection and repair of faults significantly increases long run safety, even if the fleet continues to operate after a fault has been discovered, and even if repairs may be imperfect. It follows that there is much value in monitoring, analyzing and, if warranted, correcting all non-trivial failures and their precursors.
@@
We can now see that traditional approaches to assurance, such as described for OP, give us justifiably strong confidence that the system satisfies its critical properties; this can then be augmented by testing and operational experience to deliver (via CBI and Bootstrapping) justified confidence that the system satisfies its dependability goals, and this can be reinforced by monitoring during operation. In the following sections, we consider systems with AI and ML components that do not lend themselves to traditional methods of assurance and may therefore be ineligible for CBI, and we explore how, and to what extent, dependability can nonetheless be achieved and guaranteed.
@@ 2 Assurance for Systems Extended with AI and ML
For traditional assurance, there must be good reasons why we believe the system achieves its dependability goals and those reasons are documented and justified in its assurance case. Systems that use AI and ML pose challenges to this approach because, rather than performing actions that are effective and safe for reasons that can be articulated and verified, a system that uses ML operates by learning suitable behavior during a period of training. Training typically defines empirically effective “weights” in a deep neural network (DNN); there will often be millions, or even billions, of individually adjusted weights. [^7] There are other ML techniques, such as Support Vector Machines [161] but, despite different mechanisms, these pose similar challenges to assurance as do neural networks. Similarly, the specific ML architecture in which neural networks are employed, such as reinforcement and inverse reinforcement learning, or large generative language and diffusion models, has little impact on the fundamental difficulty of assuring ML. The hope is that if the system works correctly on the training examples, then it will work correctly on all similar examples.
@@
An alternative to ML is symbolic AI, which uses automated deduction (theorem proving) to derive conclusions from premises composed of a set of axioms describing some aspect of the world plus observations about the current state of the world. It is possible to guarantee validity of some methods of automated deduction (e.g., SAT and SMT solvers with certificates [120]), but soundness depends on the choice of premises, which may be unvalidated and derived from an informal or empirical model, and also on the computational resources available (deduction generally requires exponential time, or worse, so we may need to accept whatever partial analysis can be accomplished in fixed time). “Expert systems” were a type of symbolic AI popular in the 1980s where deductive procedures were applied to a collection of “rules” that axiomatized some domain. The concern, and one of the reasons for the demise of these systems, was that individually reasonable rules could collectively be inconsistent or incomplete, resulting in faulty output [143]. Modern symbolic AI uses improved technology but its challenge to assurance remains largely unchanged. A combination of symbolic AI with ML is a popular current approach known as neurosymbolic AI. The strengths and weaknesses of the two approaches seem to be complementary, but this does not assuage their assurance problems.
@@
Traditional assurance requires good understanding of how the system works because tasks such as intent verification must show that certain properties hold in all circumstances. This is infeasible for most AI and ML components because we lack detailed understanding of their operation, [^8] Recently, there has been progress in associating learned “concepts” with specific clusters of artificial neurons [166], but this is still some way from the understanding required for assurance. but an alternative or constituent part of the overall system assurance process can be to check that properties hold in the circumstances actually encountered during operation. This is runtime verification [142] or, more boldly runtime certification [146], where components, often generically referred to as guards (or monitors [121, 71]), are added to the system to check its behavior against its required or specified properties, or conservative simplifications of these. If a check fails, then the system must take some remedial action to maintain or restore safety. Both checking and remediation add complexity to the system and may themselves introduce failures and hazards, so this approach requires careful engineering [100]. Nonetheless, a plausible approach is to guard AI and ML elements with conventionally engineered components that perform runtime verification and can be assured in the conventional way. This approach is endorsed in some industry guidelines such as F3269-17 for unmanned aircraft [7].
@@ 2.1 Runtime Verification
To investigate runtime verification for systems with AI and ML components, we need some general understanding of the likely overall system architecture and the properties that will be checked. The top-level structure of almost any system that employs AI or ML follows from a single insight, which is that any entity that interacts effectively with some aspect of the world must have a model of that aspect of the world [45]. In particular, cyber-physical systems are always based on a model of the controlled “plant” and its environment [59]. For pure control systems, the model may be a collection of differential equations; for systems that involve procedures it may add state machines; and for full CPS it will include integrated formalisms such as timed and hybrid automata. [^9] A model can be thought of as a simulation (or “digital twin”) of the plant and its environment; “model based system/software engineering” (MBSE) builds and experiments with explicit simulations of these plus the intended system using tools such as StateFlow/Simulink and derives the requirements and specification, and sometimes the implementation (and sometimes does that automatically), from the system simulation. A criticism of this approach is that the focus on simulation introduces implementation concerns too early into the process (it is difficult to simulate requirements stated as constraints unless theorem proving is used, so specifications are often substituted for requirements) and thereby compromises dependability requirements validation.
@@
The model may be used in development but be represented only indirectly in the final system (e.g., as control laws in classical control engineering), or it may be partially defined at design time and built in to the system with parameters that can be adjusted at runtime (as in adaptive control systems), or it may largely be constructed at runtime and represented explicitly within the system (as in model-predictive control and autonomous systems). A characteristic of most systems that employ AI and ML is that these capabilities are used in building and maintaining a model of the “world” (i.e., its environment, such as the locations of other road users) at runtime; this function is generally referred to as the perception (sub)system. Optionally, AI and ML may also be employed in the action (sub)system, which uses the world model to calculate and execute behavior that will advance the system’s progress toward its goal, while maintaining dependability. [^10] There is often substructure within the action system: generally a decision system that plans what to do (e.g., “move into the nearside lane between cars a and b”) and an execution system that manipulates the controls (accelerator, brakes, steering, horn) to perform that plan. These details are unimportant at our level of discussion.
@@
It is often possible to guard AI or ML-generated actions with highly assured conventional software that checks their safety against the world model. However, if the world model is constructed by a perception system that uses AI or ML then we have to ask how its own accuracy can be assured or guarded: it does no good to run safety checks against a faulty model. One possibility is for the guard to use a different model that can be assured.
@@
We can therefore distinguish two classes of guarded AI-enabled systems according to the nature of their runtime verification. Assuredly guarded systems are those whose safety and other critical properties can be checked and enforced by assured guards that do not themselves use AI or ML, neither for perception nor action. In cases where the guards are so strongly assured that they can be considered “possibly perfect” (or “probably fault-free”), it is possible to make very strong claims for dependability of the overall system [112, 183]. The second class is unassuredly guarded systems, where the guards themselves depend on AI or ML, typically in their perception systems. See, for example, the architectures listed on page 2.4, where those that include a traditionally assured backup guard are assuredly guarded and others (except number 1, which is not guarded) are unassuredly guarded.
@@
An example of the first class is an unmanned aircraft controlled by AI and ML that is deemed safe as long as it remains within some specified portion of airspace (a “geofence”). This is a constraint that can be checked by a conventional navigation system and enforced using a conventional guidance system as an override when violations are detected [48]. Many robots can be guarded by “virtual cages” of this kind. Notice also that when formulating the initial “concept of operation” for a system, it may be possible to adjust the concept so that assurable guards become feasible. For example, rather than an autonomous shuttle bus sharing its route with other vehicles and pedestrians, it could use a dedicated and isolated track, and would not then require a sophisticated AI-based perception system.
@@
An example of the second class is a self-driving car where the action guard checks AI-generated actions for safety and other dependability properties against a model of its environment—i.e., location of other road users and pedestrians, interpretation of traffic signs and markings, detection of road layout, etc. To verify the guard we need assurance for accuracy of its model, and to do this in the traditional way we would want to see its requirements (i.e., what it should represent and what it means for it to do so with acceptable accuracy), its hazards (e.g., inaccuracies that can lead to harm), together with specifications and an implementation that can be shown to achieve the requirements and mitigate the hazards. Typically, none of this is present for models constructed by an AI perception system that uses ML to interpret its sensors, although recent developments are starting to explore some of these topics [76, 39, 25].

### v3

@@ 1.3 From Assurance to Dependability
In commercial airplanes, “catastrophic failure conditions” (those “which would prevent continued safe flight and landing”) must be so unlikely that they are “not anticipated to occur during the total operational life of all airplanes of a given type” [77, Section 3.2.4]. The “total operational life of all airplanes of a given type” is about $10^{8}$ to $10^{9}$ flights for modern airplanes. With an average flight duration of about 90 minutes, this requires a critical failure rate no worse than about $10^{-9}$ per hour. [^6] There are many measures for reliability, such as failure rate, probability of failure on demand, probability of fatality per mile, etc. And the underlying system models may use discrete or continuous time, with Bernoulli or Poisson failure processes, etc. The general conclusions drawn here are robust to all these choices and so we do not describe them in detail.
@@
Cars are among the most dangerous consumer goods with about 40,000 deaths per year in the United States and a fatal accident rate of a little over 10 per billion miles. It is intended that self-driving cars should be safer than human drivers, so it might seem reasonable (even though driving error is not the only cause of accidents) to set the target at no more than 1 death per billion miles, which is around $10^{-7}$ per hour, given an average speed of 30 mph. However, that is a rather technical assessment that pays no attention to likely public reaction. Herbert Diess, the former CEO of Volkswagen is quoted on their website with a more realistic assessment: “A ratio of ten-to-one is nowhere near good enough. We have approximately 3,200 traffic fatalities in Germany each year. It would be a disaster if we had even 320 deaths due to driverless cars.” Thus, it is plausible that the safety target for self-driving cars should be 100 or even 1,000 times better than human drivers, [^7] This sensitivity is illustrated by a recent accident in San Francisco: a car with a human driver hit a pedestrian and threw them under the wheels of a Cruise self-driving taxi, which then dragged the victim for several yards. Subsequently, Cruise lost their license to operate in San Francisco (although their poor initial response to the incident probably contributed to this harsh reaction) [136]. which brings us into, or even beyond, the requirements for commercial aircraft [138]. But commercial aircraft do get assured and certified and their safety record justifies this, so why is $10^{-9}$ seen as such a challenge?
@@
The answer is that we do not assure aircraft solely by testing. If we test a system for $n$ hours and observe no failures, then in the absence of other information, the best prediction we can make is that the likelihood of no failures in another $n$ hours is about 50–50 [163, p. 73, and sidebar on p. 74]. Hence, to secure assurance for failure rates of $10^{-9}$ per hour we would need to test the system for about $10^{9}$ hours, or around 115,000 years. Even with 1,000 copies of the system on test, this is still well over 100 years of continuous operation and is completely infeasible [52, 163, 125].
It is not just the problems of developing sufficient test data that makes this assurance route challenging and often infeasible. For modest values such as $10^{-3}$ it might seem feasible to assess probability of failure directly by testing, but ensuring statistical validity is a very demanding process where the tests must be selected randomly from the operational profile [58]. It is demanding because the test harness must reproduce the actual environment, the operational profile must be accurate and must be sampled accurately (simply using the system “randomly” does not do this), the test oracle (which determines success or failure of each test) must be correct, and the number of tests must be large enough to deliver a statistically valid conclusion (and there must be no failures among them—the number of tests needed in the presence of failures is astronomical). In addition, statistical extrapolation rests on assumptions that the future will be like the past. Validating this requires analysis of all software and system components, plus the test harness and its equipment, to show there are no time dependent effects such as state accumulation, security exposure and so on.
Thus, assurance by testing alone is impossible for critical systems: the required number of tests is too large and the assumptions too onerous. Consequently, the idea behind classical methods of assurance for critical systems is to develop justified confidence that the system contains few or no faults. We do this by examining the design and construction of the system. From confidence in absence of faults we predict low probability of failure. This prediction responds to the “absence of other information” qualifier on the efficacy of testing stated earlier. As explained below, the combination of confidence in the absence of faults and a feasible amount of testing then provides system assurance.
It is fairly straightforward to reason from low likelihood of faults to low probability of failure, but assurance does not tell us the likelihood of faults: what it does is provide us with confidence that this likelihood is low. Hence we need some method for quantifying confidence that will support an inference to probability of failure.
@@
Confidence can be expressed as a subjective probability, so if we are $95\%$ confident that traditional assurance works, that means we estimate there is only 5% chance that the assured system contains faults. We can now use testing to explore the existence of those potential faults but, unlike the previous case, we know something about the system so when we see $n$ hours with no failures, we can conclude (by what is called Conservative Bayesian Inference, CBI) that we are likely to see another $10n$ with no failures [226, 252]. This reduces the amount of testing required and another idea reduces it still further. This is Bootstrapping [33]: we need assurance for $10^{9}$ hours, but this is over the lifetime of the system. When the system is first deployed, we might be satisfied to know there will be no dependability failures in the first year, and we will have only a few instances of the system operational in that time. So we might need confidence for only, say, $10^{5}$ hours, and testing for this, given prior assurance and CBI, requires only $10^{4}$ hours, which is perfectly feasible. After the first year, we might seek confidence for the next year and for the larger number of systems now installed, but we will have the operational experience of the first year and that should be sufficient (given the tenfold multiplier of CBI) to deliver the required confidence in safe operation, and so on for subsequent years. As with statistical testing, bootstrapping relies on arguing that future behavior will be like the past, so that topics such as changes in the environment and time dependent effects such as state accumulation, software updates and security exposure all need to be addressed.
Koopman criticizes bootstrapping [138, Section 8.4] when this is based solely on tests and accumulating experience, but accepts that it is sound when, as here, it is founded on justified prior confidence. He is correct, however, that bootstrapping exposes early adopters to greater risk: in the first year, we are confident the probability of failure is sufficiently low that we will not see a failure in that first year of operation, but in the second year we are confident we will not see a failure in the next year of operation (i.e., two years in total), and so the estimated probability of failure must be lower in this period than the previous one since the exposure is twice as long. This is ethically sound, however, because although the estimated probability of failure is greater in early periods, it is always sufficient to provide adequate assurance of safety. In terms of the UK ALARP approach (“As Low As Reasonably Practicable”) [197], we show the initial risk is tolerable and then use successful operating experience to further reduce our risk estimates within the ALARP region.
@@
Bootstrapping is based on failure-free operation. But suppose experience in operation does reveal a failure, hopefully not catastrophic (in commercial aircraft, for example, it is required that no catastrophic failure may be caused by a single fault). Any such failure reveals an unanticipated fault, arrival rate, circumstance, or hazard, and these could be precursors to a catastrophic failure. Consequently, commercial airplanes operate in a legal and ethical framework where all incidents and accidents are promptly reported and dispassionately investigated. The FAA issues Airworthiness Directives mandating workarounds or corrections to detected faults; in extreme cases it may temporarily ground the fleet (as it did for the 737 MAX in 2019–2020). Bishop [29] constructs a statistical model for this scenario and shows that, under plausible assumptions, detection and repair of faults significantly increases long run safety, even if the fleet continues to operate after a fault has been discovered, and even if repairs may be imperfect. It follows that there is much value in monitoring, analyzing and, if warranted, correcting all non-trivial failures and their precursors. We will see in Section 2.1 that this process can be systematized with “Safety Performance Indicators.”
@@
We can now see that traditional approaches to assurance, such as described for OP, give us justifiably strong confidence that the system satisfies its critical properties; this can then be augmented by testing and operational experience to deliver justified confidence (via CBI, Bootstrapping, and a case that addresses their assumptions) that the system satisfies its dependability goals, and this can be reinforced by monitoring during operation.
We will sometimes use expressions such as weak, modest, or strong assurance. We intend these to be interpreted as qualitative indications of the failure rates they are able to support. We have already seen that rates on the order $10^{-9}$ are required for fatal accidents in airplanes and cars, and we will say this requires strong assurance. Some safety concerns about AGI systems speak of existential risk and we might say this should require exceptional assurance such as $10^{-12}$, while modest and weak assurance might correspond to $10^{-6}$ and $10^{-3}$, respectively. These latter levels correspond to those conferred by the lower levels of established standards such as DO-178C (DALs C and D) [200], IEC 61508 (SILs 1–3) [110] etc.
All these allocations of risk and assurance are speculative: the actual values would depend on the number of systems deployed, their usage, and the nature of the harms they may cause. One of the conclusions of this report is that the risks from AI applications need further analysis to place them in the risk spectrum so that we can relate the AI assurance challenge to known approaches.
The roles of testing, analysis, and architecture change as the criticality of the system increases. At modest criticality, operational experience and statistical testing can be combined with modest analyses to justify assumptions and produce a plausible assurance case. As criticality increases, evidence from program analysis evidence becomes more dominant—e.g., using formal methods to verify absence of faults—with test and experience evidence playing a supporting role in validating assumptions (e.g., of tool efficacy). If analysis is difficult or infeasible, as it is for many AI systems, then the role of architectural mechanisms such as runtime verification becomes more important
In general, assurance needs analysis and architectural mechanisms combined with testing. The combinations can be analyzed and justified using CBI for high assurance, a related theory known as N/T (“N over T”) [31, 30] for modest levels, and “probability of perfection” [162] for architectural mechanisms.
In the following sections, we consider systems with AI and ML components that do not lend themselves to traditional methods of assurance and we explore how, and to what extent, dependability can nonetheless be achieved and guaranteed.
@@ 2 Assurance for Systems Extended with AI and ML
For traditional assurance, there must be good reasons why we believe the system achieves its dependability goals and those reasons are documented and justified in its assurance case. Systems that use AI and ML pose challenges to this approach. For example, symbolic AI, which is sometimes seen as an alternative to ML, uses automated deduction (theorem proving) to derive conclusions from premises composed of a set of axioms describing some aspect of the world plus observations about the current state of the world. It is possible to guarantee validity of some methods of automated deduction (e.g., SAT and SMT solvers with certificates [172]), but soundness depends on the choice of premises, which may be unvalidated and derived from an informal or empirical model, and also on the computational resources available (deduction generally requires exponential time, or worse, so we may need to accept whatever partial analysis can be accomplished in fixed time). “Expert systems” were a type of symbolic AI popular in the 1980s where deductive procedures were applied to a collection of “rules” that axiomatized some domain. The concern, and one of the reasons for the demise of these systems, was that individually reasonable rules could collectively be inconsistent or incomplete, resulting in faulty output [202]. Modern symbolic AI uses improved technology but its challenge to assurance remains largely unchanged.
In a similar challenge to assurance, rather than performing actions that are effective and safe for reasons that can be articulated and verified, ML components operate by learning suitable behavior during a period of training. Training typically defines empirically effective “weights” in a deep neural network (DNN); there will often be millions, or even billions, of individually adjusted weights. [^8] There are other ML techniques, such as Support Vector Machines [224] but, despite different mechanisms, these pose similar challenges to assurance as do neural networks. Similarly, the specific ML architecture in which neural networks are employed, such as reinforcement and inverse reinforcement learning, or large generative language and diffusion models, has little impact on the fundamental difficulty of assuring ML. The hope is that if the system works correctly on the training examples, then it will work correctly on all similar examples, but there are no strong methods to guarantee this.
A combination of symbolic AI with ML is a popular current approach known as neurosymbolic AI. The strengths and weaknesses of the two approaches seem to be complementary, but this does not assuage their assurance problems.
@@
Traditional assurance requires good understanding of how the system works because tasks such as intent verification must show that certain properties hold in all circumstances. This is infeasible for most AI and ML components because we lack detailed understanding of their operation [^9] Recently, there has been progress in associating learned “concepts” with specific clusters of artificial neurons [230], but this is still some way from the understanding required for assurance. and without this, testing is inadequate. However, an alternative or constituent part of the overall system assurance process can be to check that properties hold in the circumstances actually encountered during operation. This is runtime verification [201] or, more boldly runtime certification [205], where components, often generically referred to as guards (or monitors [173, 99]), are added to the system to check its behavior against its required or specified safety properties, or conservative simplifications of these. If a check fails, then the system must take some remedial action to maintain or restore safety. Runtime checking is not always possible (e.g., for perception, where we have no independent knowledge of the world) and both checking and remediation add complexity to the system and may themselves introduce failures and hazards, so this approach requires careful engineering [142]. Nonetheless, a plausible approach is to guard suitable AI and ML elements with conventionally engineered components that perform runtime verification and can be assured in the conventional way. This approach is endorsed in some industry guidelines such as F3269-17 for unmanned aircraft [8]. To be coherent, the overall assurance case should include claims that are supported by subcases for the runtime checking and remediation components and an argument that integrates these with the case for the primary system.
2.1 Safety Performance Indicators
Before examining the details of runtime verification, we discuss a valuable secondary use of the same technology. Runtime verification checks critical claims in the overall assurance case at runtime and intervenes if these are violated. We can imagine checking lesser claims that appear in subcases supporting critical claims as indicators of potential violations of critical claims. For example, a critical claim might be that a self-driving car will not collide with pedestrians. A lesser, supporting claim might be that the car should provide a one meter buffer between itself and pedestrians. Violations of the lesser claim suggest that things are not working as intended and that in-service adjustments may be needed in the system and its assurance case.
Runtime checks of lesser claims are called Safety Performance Indicators (SPIs); Johansson and Koopman [120] define them as data-supported metrics that provide a threshold on the validity of assurance claims. SPIs may be collected and evaluated per system (so excessive violations of the one meter buffer may indicate that this car’s sensors are faulty or dirty) or over the whole population (so statistically excessive violations of the one meter buffer may indicate that the relevant part of the system and/or its safety case are deficient). In some cases, the system assurance case may have “dynamic” elements that reference SPIs and adjust confidence in assurance at runtime [68]. We do not consider this at odds with the requirement that assurance cases should be indefeasible: indefeasibility applies to the critical overall claims, while SPIs measure thresholds on lesser claims. If concern propagates from lesser to critical claims, then it means that the assurance case is deficient, not that the criterion for indefeasible soundness is flawed. In addition to SPIs that function as leading indicators, warning of possible problems ahead, lagging indicators can also be useful by providing feedback on how well things are going. The UL4600 standard for evaluation of autonomous products has extensive requirements for SPIs [235, Section 16][139].
@@ 2.2 Runtime Verification
To investigate runtime verification for systems with AI and ML components, we need some general understanding of the likely overall system architecture and the properties that will be checked. The top-level structure of almost any system that employs AI or ML follows from a single insight, which is that any entity that interacts effectively with some aspect of the world must have a model of that aspect of the world [63]. In particular, cyber-physical systems are always based on a model of the controlled “plant” and its environment [82]. The model can be thought of as a simulation of relevant aspects of the world that allows the system to predict its behavior and thereby choose appropriate actions.
We find it useful to distinguish two aspects of world models. A local model represents the configuration of the system’s “world”: that is, the state of all relevant objects and attributes in its immediate environment, in sufficient detail and precision for the system to perform its task. The local model is typically built by the system’s sensors and associated subsystems. Additionally, there is a domain model that represents how the elements in the local model interact: that is, how the world “works.” When we say “world model” or just “model” we mean the combination of local and domain model.
For pure control systems, the domain model may be a collection of differential equations and the local model will be the sampled values of the variables over which it is defined. For systems that involve procedures the domain model may include state machines, and for CPS it may be described by integrated formalisms such as timed and hybrid automata.101010“Model based system/software engineering” (MBSE) builds and experiments with explicit simulation models using tools such as StateFlow/Simulink and derives the system requirements and specification, and sometimes its implementation (and sometimes does that automatically), from the simulation. A criticism of this approach is that the focus on simulation introduces implementation concerns too early into the process (it is difficult to simulate requirements stated as constraints unless theorem proving is used, so specifications are often substituted for requirements) and thereby compromises dependability requirements validation. For autonomous systems, we take the example of a self-driving car, where the local model concerns the immediate road layout with its obstructions, traffic signs and signals, the locations of other road users, pedestrians, and so on, and their inferred intent. The domain model describes how such objects move (or do not move), including physics for speed, acceleration, braking, momentum, centripetal force, tire friction, human reaction time and so on. The domain model is used to predict how the local model will evolve in time. In a self-driving car, it may be programmed by human designers but it can also be developed by ML. In later sections we will consider systems that infer their own domain models or are constructed on foundations that have learned them in training. In those advanced systems, the domain model may include human behavior and even societal topics such as ethics.
@@
Systems that employ AI and ML often use them in their perception (sub)system, which is used to build and maintain their local world model. Optionally, AI and ML may also be employed in the action (sub)system, which uses both local and domain models to calculate and execute behavior that will advance the system’s progress toward its goal, while maintaining dependability. [^11] There is often substructure within the action system: generally a decision system that plans what to do (e.g., “move into the nearside lane between cars a and b”) and an execution system that manipulates the controls (accelerator, brakes, steering, horn) to perform that plan. These details are unimportant at our level of discussion.
@@
It is often possible to guard AI or ML-generated actions with highly assured conventional software that uses an explicit human-developed domain model to check their safety against the local model. However, if the local model is constructed by a perception system that uses AI or ML then we have to ask how its own accuracy can be assured or guarded: it does no good to run safety checks against a faulty model. However, in some cases it is possible for the guard to use a local model that can be assured.
@@
We can therefore distinguish two classes of guarded AI-enabled systems according to the nature of their runtime verification. Assuredly guarded systems are those whose safety and other critical properties can be checked and enforced by assured guards that do not themselves use AI or ML, neither for perception nor action. In cases where the guards are so strongly assured that they can be considered “possibly perfect” (or “probably fault-free”), it is possible to make very strong claims for dependability of the overall system [162, 253]. The second class is unassuredly guarded systems, where the guards themselves depend on AI or ML, typically in their perception systems. See, for example, the architectures listed on page 2.5, where those that include a traditionally assured backup guard are assuredly guarded and others (except number 1, which is not guarded) are unassuredly guarded.
@@
An example of the first class is an unmanned aircraft controlled by AI and ML that is deemed safe as long as it remains within some specified portion of airspace (a “geofence”). This is a constraint that can be checked by a conventional navigation system and enforced using a conventional guidance system as an override when violations are detected [69]. Here, the navigation system provides a local model for the guard and its guidance override system uses a human-constructed domain model. Many robots can be guarded by “virtual cages” of this kind. Notice also that when formulating the initial “concept of operation” for a system, it may be possible to adjust the concept so that assurable guards become feasible. For example, rather than an autonomous shuttle bus sharing its route with other vehicles and pedestrians, it could use a dedicated and isolated track, and would not then require a sophisticated AI-based perception system.
@@
An example of the second class is a self-driving car where the action guard uses an assurable human-provided domain model to check AI-generated actions for safety and other dependability properties against a local model of its environment—i.e., location of other road users and pedestrians, interpretation of traffic signs and markings, detection of road layout, etc. But to fully verify the guard we need assurance for accuracy of its local model, and to do this in the traditional way we would want to see a strong argument that the model satisfies its dependability requirements and that these are sufficient to mitigate its hazards. As we examine in the next section, this is typically infeasible for models constructed by an AI perception system that uses ML to interpret its sensors (although recent developments are starting to explore some of these topics [106, 55, 36]).